he all-equity firm Metallica Heavy Metal Mining (MHMM) Corporation wants to diversify its operations. Some recent financial information for the company is shown here:

  Stock price $ 69
  Number of shares 30,000
  Total assets $ 8,700,000
  Total liabilities $ 3,600,000
  Net income $ 600,000

MHMM is considering an investment that has the same PE ratio as the firm. The cost of the investment is $640,000, and it will be financed with a new equity issue. The return on the investment will equal MHMM’s current ROE.

c. What will happen to the market-to-book ratio? (Do not round intermediate calculations and round your answers to 4 decimal places, e.g., 32.1616.)
d. What will happen to the EPS? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
e. What is the NPV of this investment?
Answer #1

c)

Market to Book ratio = (Market value of shares/ Book Value of shares)

Market value of shares = 69 * 30,000 = 2,070,000

Book value of shares = Book value of assets - BV of equity

BV of assets = 8,700,000

BV of liabilities = 3,600,000

BV of shares = 8,700,000 - 3,600,000 = 5,100,000

Market to Book ratio = 2,070,000/5,100,000 = 0.4059

d)

EPS = Earnings per share = (Net Income / Number of shares) = 600,000 / 30,000 = 20

e)

ROE on the investment = Net Income / Equity = 600,000/ 5,100,000 = 11.76%

Initial investment = 640,000

PV of cash flows = (1+ROE) * 640,000 = 715,294

NPV of cash flows = 715,294,117 - 640,000 = $ 75,294